# The HITRAN2024 methane update

### v2

@@ 2.4.2 HITRAN update in the 3900–4900 cm-1 region
Spectra $S1$ and from Birk et al. [birk_esa_2017] ($S3$), including CRDS spectra not shown here, were used for the octad region. For the same reason that $S1$ shows a lower SNR, $S3$ was used for its window of 3900–4500 cm-1 instead.
Figure 3: Residuals obtained at 296 K and 0.17143 hPa between the experiment $S3$ from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 4: Residuals obtained at 198 K and 0.16941 hPa between the experiment $S3$ from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
@@
Unlike in the pentad region, several new line lists presented a clear improvement over large portions of the spectrum, as illustrated in Figures 3 and 4. This result alleviated the need for line-by-line analysis, with a few exceptions that were then fixed. In these large portions, the new lists were merged with HITRAN2020, adding new lines and updating existing ones unless no significant changes were made by conditions similar to those detailed in Sec. 2.2.2 , depending on the precision of the parameters. The residuals were then analyzed, keeping HITRAN2020 parameters if needed, ensuring that only the best parameters remained. This method was used for the data from Rodina et al. [rodina_improved_2021], between 4100 and 4190 cm-1, improving both CH4, 13CH4, and CH3D. Their line list extended to 4300 cm-1, but that region was previously updated in HITRAN2020 for TROPOMI [hu_operational_2016]. A list provided by Sung et al. for 13CH4 has been added as well between 4000–4700 cm-1, including the region between 4190–4300 cm-1 which replaces some lines, or intensities, incorrectly assigned to the main isotopologues. The list from Nikitin et al. [nikitin_preliminary_2013] improving CH3D between 4000–4554 cm-1 has been added, adding the lines missing in HITRAN2016 [gordon_hitran2016_2017] when the preliminary version of their data was first used. Finally, the empirical list from Toon (atm.161 [toon_atmospheric_2022]) replaced the region between 4341–4900 cm-1 for lines above the intensity of $10^{-27}$ cm//molecule, improving CH4, 13CH4, and CH3D.
We observed an approximately systematic shift in the transition wavenumber of lines from Daumont et al. [daumont_new_2013]. Applying a constant shift of 0.001 cm-1 appeared to improve the residuals, particularly below 4200 cm-1. The effect is limited above 4375 cm-1 as the addition of atm.161 list already improved that region appreciably. The final residuals for the latest HITRAN versions are shown in Figure 5, confirming the great improvement above 4375 cm-1, after keeping the region unchanged in the last 3 versions.
@@ 2.5.1 IAO analyses of JPL FTS measurements of 13CH4
Figure 6: An overview of the 13CH4 tetradecad line lists available in HITRAN2024. Newly added line parameters for 13CH4 are highlighted in magenta. The TheoReTS ab initio based variational list [rey_new_2018] with intensity cut-off $1\times 10^{-29}$ cm/molecule (grey dots) is plotted as a background overlapping with the colored data
Starikova et al. [starikova_assignment_2024] reported the first assignments of 13CH4 absorption bands belonging to the lower part of the tetradecad. This work focused on the analysis of a high-resolution Fourier-transform infrared spectrum of 13C-enriched methane in the 4970–5300 cm-1 range (Fig. 6). The spectrum was recorded over a broad range (4970–6200 cm-1) at 298 K using a 20.94 m path cryogenic Herriott cell vacuum-coupled to a Fourier transform spectrometer, Bruker IFS-125HR, at the Jet Propulsion Laboratory. The retrievals of line intensities from the experimental spectrum were carried out using the SpectraPlot [nikitin_visualization_2011] program, with the Voigt line profile at fixed values $\gamma_{\mathrm{self}}=0.08\ \mathrm{cm}^{-1}\mathrm{atm}^{-1}$, $n_{\mathrm{self}}=0.85$, and $\delta_{\mathrm{self}}=-0.018\ \mathrm{cm}^{-1}\mathrm{atm}^{-1}$. The uncertainties in the intensities of isolated lines were estimated to be about 10–15%, while those of blended lines were about 20–40%. Initial spectral assignments were obtained using the effective Hamiltonian model and effective dipole transition moment parameters, both derived from the previous work on cold 13CH4 spectra recorded at 80 K [starikova_13ch4_2018, starikova_assignment_2019, starikova_assignment_2016]. The published experimental list contains 1642 lines, of which 1600 transitions were assigned to four vibrational bands $4\nu_{4}$, $\nu_{2}+3\nu_{4}$, $\nu_{3}+2\nu_{4}$ and $\nu_{1}+2\nu_{4}$ of the tetradecad system, including 22 sub-bands with rotational quantum numbers up to $J_{\mathrm{max}}=16$. The parameters of the effective Hamiltonian and the dipole transition moment allowed the description of 1548 fitted experimental line positions with an rms deviation of $1.7\times 10^{-3}\mathrm{cm}^{-1}$, and the reproduction of a set of 674 measured line intensities with an rms deviation of 6.8%.
Figure 7: Comparison between the experimental spectrum of 13CH4 at 298 K (lower panel) with simulation based on the effective Hamiltonian and dipole transition moment models (middle panel) and ab initio TheoReTS [rey_new_2018] calculations (upper panel) in the 5000–5220 cm-1 tetradecad region. The blown-up part at the bottom corresponds to the narrow spectral interval around 5178 cm-1.
A comparison of the observed spectrum with the simulations using calculated effective parameters and the TheoReTS [rey_new_2018] line list, depicted in Fig. 7, shows the excellent quality of the calculations for this molecule as well as the improvements due to the reduction of experimental data.
@@ 2.5.2 HITRAN update in the 4970–5300 cm-1 region
The new data from Starikova et al. [starikova_assignment_2024] partially fills the gap of 13CH4 in the lower part of the tetradecad region. More than 1600 new lines were incorporated in HITRAN2024 (see Fig. 6). No 13CH4 lines were present in this region in HITRAN2020.
@@ 2.6.1 NIST, DLR & USTC measurements
The tetradecad spectral region has been the subject of extensive study, including the WKLMC empirical line list [campargue_2013], the GOSAT 2014 line list [nikitin_gosat-2014_2015] (which was the basis of both HITRAN2016 and HITRAN2020 for this spectral region), the widely used empirical line list of Devi 2015 [devi_self-_2015], and the Nikitin 2017 analysis of WKLMC [nikitin_analysis_2017]. Significant differences exist between these line lists, complicated by correlation between parameters in these dense and complex spectra and the difficulty in identifying and assigning weaker transitions. These complications lead to line lists that trade parameter accuracy for self-consistency, thus impeding efforts to refine line parameters.
Recently, this problem has been examined in the methane tetradecad via three separate sets of experiments focused on the strong $2\nu_{3}$ band of 12CH4, which represents $\approx{67\%}$ of the total summed intensity of the R-branch spectral region from 6010 cm-1 to 6120 cm-1. Each experiment was tailored for optimal determination of particular spectroscopic parameters: saturation spectroscopy for transition frequencies, linear absorption spectroscopy in the Doppler-broadened domain for absolute transition intensities and ratios, and line shape and line mixing parameters determined in the pressure broadened domain. Transition frequencies for 53 transitions in the $2\nu_{3}$ R-branch were determined with kHz-level uncertainty through saturation spectroscopy by Votava et al. [votava_2022]. These highly accurate values were subsequently used as constraints in the determination of molecular line intensities in the Doppler-broadened domain through independent experiments employing CRDS at NIST and FTS at DLR as reported by Reed et al. [reed_multilaboratory_2025], which additionally replaced $\approx 2600$ weaker non-$2\nu_{3}$ lines from the Nikitin 2017 line list. Permille-level uncertainties were demonstrated (evaluated by relative agreement between the two independent measurements) for 64 strong transitions, covering the $2\nu_{3}$ R-branch from $m=$ (1–11). These line intensities were employed by Yin et al.[yin_2025] in further CRDS studies involving elevated pressure of CH4 broadened by N2 to determine line shape and line mixing parameters.
@@
Agreement with TheoReTS [rey_theorets_2016] is excellent (at the several permille level) for the assigned $2\nu_{3}$ transition intensities, however, the lack of assignments in the experimental line list for many weaker (non-$2\nu_{3}$) transitions makes direct comparison difficult. Limited experimental signal-to-noise or fitting correlation requires the inclusion of many unassigned weak transitions in the experimental linelist to fully model the observed spectra. These transitions likely represent the sum of a number of weaker transitions predicted by THeoReTS.
@@ 2.6.2 LIPhy DAS measurements of 12CH3D
In spite of the small magnitude of the CH3D/CH4 amount ratio, (nominally $5\times 10^{-4}$ at natural abundance), CH3D has a large relative contribution to the methane absorption in the region (up to 70 % near 6300 cm-1 at 80 K) [wang_high_2010]) and contributes greatly to the absorption in the 1.58 µm methane transparency window [lu_ch3d_2011]. Nevertheless, the $3\nu_{2}$ band is not isolated but superimposed on a congested spectrum of lines involving other weaker bands which have to be characterized for a proper description of the absorption in the region. As described in [brown_methane_2013], since its 2012 edition, the HITRAN list was extended to the 6204-6510 cm-1 region by including for the 6204–6394 cm-1 interval, the empirical line list constructed in Lu et al. [lu_ch3d_2011] using diode laser absorption spectroscopy (DAS) of highly enriched CH3D at 296 K. This list of about 5500 lines included about 175 rovibrational assignments and more than 2700 empirical values of the lower state energy, $E^{\prime\prime}_{\mathrm{emp}}$, obtained from the ratios of intensities measured at 296 K and 80 K (2T-method).
The update of CH3D in this spectral region is based on the recent work of Ben Fathallah et al. [benfathallah_ch3d_2024] which extends that of Lu et al. The region of the DAS recordings is increased to include the 6099–6530 cm-1 region (11189 lines) and about 4800 $E^{\prime\prime}_{\mathrm{emp}}$ values were obtained by the 2T-method. Relying on the position and intensity agreements with the TheoReTS variational line list [rey_accurate_2014, rey_theorets_2016] about 2890 transitions could be fully rovibrationally assigned to twenty bands, fifteen of them being newly reported. All the reported assignments were confirmed by Ground State Combination Difference (GSCD) relations i.e. all the upper state energies (about 1370 in total) have coinciding determinations through several transitions (up to 8). The intensity sum of the transitions assigned between 6190 and 6530 cm-1 represents 76.9 % of the total experimental intensities at 296 K.
@@ 2.6.3 HITRAN update in the 5900–6530 cm-1 region
Figure 9: Residuals obtained at 294 K and 0.5 hPa between the experiment from Ref. [birk_measurement_2025] and Voigt calculations from the line lists used for the HITRAN2024 update.
The DLR/NIST line list [reed_multilaboratory_2025, yin_2025] based on Nikitin 2017 [nikitin_analysis_2017] replaced the GOSAT 2014 12CH4 transition wavenumbers, and intensities in their region, for more than 2500 transitions. In several manifolds, better residuals were achieved with fewer lines than previously included in HITRAN2020 (Fig. 9). As such, some lines were removed to accommodate for these changes as well. The new pressure-induced parameters were used to replace the previous values, as discussed below in Sec. 3.
The new list from Ben Fathallah et al. [benfathallah_ch3d_2024] that uses their experimental 12CH3D, and variational calculations if not available, was included in HITRAN2024. More than 5800 new transitions supplement the lower wavenumber range of the previous data. These data comprise 4800 lines providing updates on the positions, intensities, and energy levels, as well as new assignments.
@@ 2.7.1 PLDL Dual frequency comb
Malarich et al. [malarich_dual_2021] present new broadband, high-resolution dual frequency comb spectroscopy (DCS) of pure methane up to 1000 K and 300 Torr in the methane icosad (6770–7570 cm-1). The dual comb spectrometer spans 870 cm-1 with 0.00667 cm-1 point spacing and a wavenumber uncertainty of 10-5 cm-1. Light from the spectrometer was passed twice through a 46-cm-long optical cell (92 cm total pathlength) placed in a uniform three-zone furnace.
Thirteen total spectra were analyzed at five temperatures from 296-1000 K and 18-300 Torr. Simulations in the icosad region with HITRAN2020 show good agreement with the measured spectra at room temperature, but exhibit significant mismatches at higher temperatures. This is not surprising given that the empirically based HITRAN2020 icosad line list originated from spectra measured below 280 K (WKLMC [campargue_2013]). The new high-temperature spectra were used to update many HITRAN2020 lower-state energies, line positions and self shifts, self widths, and power-law temperature dependence exponents for self width and shift.
Much of the high-temperature discrepancy between HITRAN2020 and the DCS data was due to methane absorption features that had been given a default value of lower-state energy (such as $E"\textsubscript{HITRAN}=999$ cm-1). Using multispectrum fitting to the high-temperature spectra, 4283 lower-state energies are assigned, which had previously been given this default value. Existing lower-state energies were also updated for 92 lines, and the database adds 293 high-temperature lines that improve the database performance above 300 K. The method for these updates is described in detail in [malarich_dual_2021]. Fig.8 in Ref. [malarich_dual_2021] shows the distribution of new lower-state energy values.
@@
Figure 10: Representative high-temperature DCS spectrum, at 623 Kelvin, 300 Torr (brown crosses) [malarich_dual_2021]. Much of the HITRAN2020 residual (green dotted traces) is due to features that have the default $E"\textsubscript{HITRAN2020}=999.99$ cm-1. The net negative residual indicates that $E"\textsubscript{HITRAN2020}$ is too large. Orange trace shows final model with updates described here and in [malarich_dual_2021].

### v3

@@ 2.4.2 HITRAN update in the 3900–4900 cm-1 region
Spectra $S1$ and from Birk et al. [7] ($S3$), including CRDS spectra not shown here, were used for the octad region. For the same reason that $S1$ shows a lower SNR, $S3$ was used for its window of 3900–4500 cm-1 instead.
Figure 3: Residuals obtained at 296 K and 0.17143 hPa between the experiment $S3$ from Ref. [7] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 4: Residuals obtained at 198 K and 0.16941 hPa between the experiment $S3$ from Ref. [7] and Voigt calculations from the line lists used for the HITRAN2024 update.
@@
Unlike in the pentad region, several new line lists presented a clear improvement over large portions of the spectrum, as illustrated in Figures 3 and 4. This result alleviated the need for line-by-line analysis, with a few exceptions that were then fixed. In these large portions, the new lists were merged with HITRAN2020, adding new lines and updating existing ones unless no significant changes were made by conditions similar to those detailed in Sec. 2.2.2 , depending on the precision of the parameters. The residuals were then analyzed, keeping HITRAN2020 parameters if needed, ensuring that only the best parameters remained. This method was used for the data from Rodina et al. [95], between 4100 and 4190 cm-1, improving both CH4, 13CH4, and CH3D. Their line list extended to 4300 cm-1, but that region was previously updated in HITRAN2020 for TROPOMI [48]. A list provided by Sung et al. for 13CH4 has been added as well between 4000–4700 cm-1, including the region between 4190–4300 cm-1 which replaces some lines, or intensities, incorrectly assigned to the main isotopologues. The list from Nikitin et al. [76] improving CH3D between 4000–4554 cm-1 has been added, adding the lines missing in HITRAN2016 [38] when the preliminary version of their data was first used. Finally, the empirical list from Toon (atm.161 [109]) replaced the region between 4341–4900 cm-1 for lines above the intensity of $10^{-27}$ cm//molecule, improving CH4, 13CH4, and CH3D.
We observed an approximately systematic shift in the transition wavenumber of lines from Daumont et al. [19]. Applying a constant shift of 0.001 cm-1 appeared to improve the residuals, particularly below 4200 cm-1. The effect is limited above 4375 cm-1 as the addition of atm.161 list already improved that region appreciably. The final residuals for the latest HITRAN versions are shown in Figure 5, confirming the great improvement above 4375 cm-1, after keeping the region unchanged in the last 3 versions.
@@ 2.5.1 IAO analyses of JPL FTS measurements of 13CH4
Figure 6: An overview of the 13CH4 tetradecad line lists available in HITRAN2024. Newly added line parameters for 13CH4 are highlighted in magenta. The TheoReTS ab initio based variational list [84] with intensity cut-off $1\times 10^{-29}$ cm/molecule (grey dots) is plotted as a background overlapping with the colored data
Starikova et al. [100] reported the first assignments of 13CH4 absorption bands belonging to the lower part of the tetradecad. This work focused on the analysis of a high-resolution Fourier-transform infrared spectrum of 13C-enriched methane in the 4970–5300 cm-1 range (Fig. 6). The spectrum was recorded over a broad range (4970–6200 cm-1) at 298 K using a 20.94 m path cryogenic Herriott cell vacuum-coupled to a Fourier transform spectrometer, Bruker IFS-125HR, at the Jet Propulsion Laboratory. The retrievals of line intensities from the experimental spectrum were carried out using the SpectraPlot [77] program, with the Voigt line profile at fixed values $\gamma_{\mathrm{self}}=0.08\ \mathrm{cm}^{-1}\mathrm{atm}^{-1}$, $n_{\mathrm{self}}=0.85$, and $\delta_{\mathrm{self}}=-0.018\ \mathrm{cm}^{-1}\mathrm{atm}^{-1}$. The uncertainties in the intensities of isolated lines were estimated to be about 10–15%, while those of blended lines were about 20–40%. Initial spectral assignments were obtained using the effective Hamiltonian model and effective dipole transition moment parameters, both derived from the previous work on cold 13CH4 spectra recorded at 80 K [99, 101, 98]. The published experimental list contains 1642 lines, of which 1600 transitions were assigned to four vibrational bands $4\nu_{4}$, $\nu_{2}+3\nu_{4}$, $\nu_{3}+2\nu_{4}$ and $\nu_{1}+2\nu_{4}$ of the tetradecad system, including 22 sub-bands with rotational quantum numbers up to $J_{\mathrm{max}}=16$. The parameters of the effective Hamiltonian and the dipole transition moment allowed the description of 1548 fitted experimental line positions with an rms deviation of $1.7\times 10^{-3}\mathrm{cm}^{-1}$, and the reproduction of a set of 674 measured line intensities with an rms deviation of 6.8%.
Figure 7: Comparison between the experimental spectrum of 13CH4 at 298 K (lower panel) with simulation based on the effective Hamiltonian and dipole transition moment models (middle panel) and ab initio TheoReTS [84] calculations (upper panel) in the 5000–5220 cm-1 tetradecad region. The blown-up part at the bottom corresponds to the narrow spectral interval around 5178 cm-1.
A comparison of the observed spectrum with the simulations using calculated effective parameters and the TheoReTS [84] line list, depicted in Fig. 7, shows the excellent quality of the calculations for this molecule as well as the improvements due to the reduction of experimental data.
@@ 2.5.2 HITRAN update in the 4970–5300 cm-1 region
The new data from Starikova et al. [100] partially fills the gap of 13CH4 in the lower part of the tetradecad region. More than 1600 new lines were incorporated in HITRAN2024 (see Fig. 6). No 13CH4 lines were present in this region in HITRAN2020.
@@ 2.6.1 NIST, DLR & USTC measurements
The tetradecad spectral region has been the subject of extensive study, including the WKLMC empirical line list [13], the GOSAT 2014 line list [69] (which was the basis of both HITRAN2016 and HITRAN2020 for this spectral region), the widely used empirical line list of Devi 2015 [21], and the Nikitin 2017 analysis of WKLMC [68]. Significant differences exist between these line lists, complicated by correlation between parameters in these dense and complex spectra and the difficulty in identifying and assigning weaker transitions. These complications lead to line lists that trade parameter accuracy for self-consistency, thus impeding efforts to refine line parameters.
Recently, this problem has been examined in the methane tetradecad via three separate sets of experiments focused on the strong $2\nu_{3}$ band of 12CH4, which represents $\approx{67\%}$ of the total summed intensity of the R-branch spectral region from 6010 cm-1 to 6120 cm-1. Each experiment was tailored for optimal determination of particular spectroscopic parameters: saturation spectroscopy for transition frequencies, linear absorption spectroscopy in the Doppler-broadened domain for absolute transition intensities and ratios, and line shape and line mixing parameters determined in the pressure broadened domain. Transition frequencies for 53 transitions in the $2\nu_{3}$ R-branch were determined with kHz-level uncertainty through saturation spectroscopy by Votava et al. [119]. These highly accurate values were subsequently used as constraints in the determination of molecular line intensities in the Doppler-broadened domain through independent experiments employing CRDS at NIST and FTS at DLR as reported by Reed et al. [82], which additionally replaced $\approx 2600$ weaker non-$2\nu_{3}$ lines from the Nikitin 2017 line list. Permille-level uncertainties were demonstrated (evaluated by relative agreement between the two independent measurements) for 64 strong transitions, covering the $2\nu_{3}$ R-branch from $m=$ (1–11). These line intensities were employed by Yin et al.[122] in further CRDS studies involving elevated pressure of CH4 broadened by N2 to determine line shape and line mixing parameters.
@@
Agreement with TheoReTS [83] is excellent (at the several permille level) for the assigned $2\nu_{3}$ transition intensities, however, the lack of assignments in the experimental line list for many weaker (non-$2\nu_{3}$) transitions makes direct comparison difficult. Limited experimental signal-to-noise or fitting correlation requires the inclusion of many unassigned weak transitions in the experimental linelist to fully model the observed spectra. These transitions likely represent the sum of a number of weaker transitions predicted by THeoReTS.
@@ 2.6.2 LIPhy DAS measurements of 12CH3D
In spite of the small magnitude of the CH3D/CH4 amount ratio, (nominally $5\times 10^{-4}$ at natural abundance), CH3D has a large relative contribution to the methane absorption in the region (up to 70 % near 6300 cm-1 at 80 K) [120]) and contributes greatly to the absorption in the 1.58 µm methane transparency window [56]. Nevertheless, the $3\nu_{2}$ band is not isolated but superimposed on a congested spectrum of lines involving other weaker bands which have to be characterized for a proper description of the absorption in the region. As described in [11], since its 2012 edition, the HITRAN list was extended to the 6204-6510 cm-1 region by including for the 6204–6394 cm-1 interval, the empirical line list constructed in Lu et al. [56] using diode laser absorption spectroscopy (DAS) of highly enriched CH3D at 296 K. This list of about 5500 lines included about 175 rovibrational assignments and more than 2700 empirical values of the lower state energy, $E^{\prime\prime}_{\mathrm{emp}}$, obtained from the ratios of intensities measured at 296 K and 80 K (2T-method).
The update of CH3D in this spectral region is based on the recent work of Ben Fathallah et al. [5] which extends that of Lu et al. The region of the DAS recordings is increased to include the 6099–6530 cm-1 region (11189 lines) and about 4800 $E^{\prime\prime}_{\mathrm{emp}}$ values were obtained by the 2T-method. Relying on the position and intensity agreements with the TheoReTS variational line list [87, 83] about 2890 transitions could be fully rovibrationally assigned to twenty bands, fifteen of them being newly reported. All the reported assignments were confirmed by Ground State Combination Difference (GSCD) relations i.e. all the upper state energies (about 1370 in total) have coinciding determinations through several transitions (up to 8). The intensity sum of the transitions assigned between 6190 and 6530 cm-1 represents 76.9 % of the total experimental intensities at 296 K.
@@ 2.6.3 HITRAN update in the 5900–6530 cm-1 region
Figure 9: Residuals obtained at 294 K and 0.5 hPa between the experiment from Ref. [8] and Voigt calculations from the line lists used for the HITRAN2024 update.
The DLR/NIST line list [82, 122] based on Nikitin 2017 [68] replaced the GOSAT 2014 12CH4 transition wavenumbers, and intensities in their region, for more than 2500 transitions. In several manifolds, better residuals were achieved with fewer lines than previously included in HITRAN2020 (Fig. 9). As such, some lines were removed to accommodate for these changes as well. The new pressure-induced parameters were used to replace the previous values, as discussed below in Sec. 3.
The new list from Ben Fathallah et al. [5] that uses their experimental 12CH3D, and variational calculations if not available, was included in HITRAN2024. More than 5800 new transitions supplement the lower wavenumber range of the previous data. These data comprise 4800 lines providing updates on the positions, intensities, and energy levels, as well as new assignments.
@@ 2.7.1 PLDL Dual frequency comb
Malarich et al. [60] present new broadband, high-resolution dual frequency comb spectroscopy (DCS) of pure methane up to 1000 K and 300 Torr in the methane icosad (6770–7570 cm-1). The dual comb spectrometer spans 870 cm-1 with 0.00667 cm-1 point spacing and a wavenumber uncertainty of 10-5 cm-1. Light from the spectrometer was passed twice through a 46-cm-long optical cell (92 cm total pathlength) placed in a uniform three-zone furnace.
Thirteen total spectra were analyzed at five temperatures from 296-1000 K and 18-300 Torr. Simulations in the icosad region with HITRAN2020 show good agreement with the measured spectra at room temperature, but exhibit significant mismatches at higher temperatures. This is not surprising given that the empirically based HITRAN2020 icosad line list originated from spectra measured below 280 K (WKLMC [13]). The new high-temperature spectra were used to update many HITRAN2020 lower-state energies, line positions and self shifts, self widths, and power-law temperature dependence exponents for self width and shift.
Much of the high-temperature discrepancy between HITRAN2020 and the DCS data was due to methane absorption features that had been given a default value of lower-state energy (such as $E"\textsubscript{HITRAN}=999$ cm-1). Using multispectrum fitting to the high-temperature spectra, 4283 lower-state energies are assigned, which had previously been given this default value. Existing lower-state energies were also updated for 92 lines, and the database adds 293 high-temperature lines that improve the database performance above 300 K. The method for these updates is described in detail in [60]. Fig.8 in Ref. [60] shows the distribution of new lower-state energy values.
@@
Figure 10: Representative high-temperature DCS spectrum, at 623 Kelvin, 300 Torr (brown crosses) [60]. Much of the HITRAN2020 residual (green dotted traces) is due to features that have the default $E"\textsubscript{HITRAN2020}=999.99$ cm-1. The net negative residual indicates that $E"\textsubscript{HITRAN2020}$ is too large. Orange trace shows final model with updates described here and in [60].